# What no one has seen before: gravitational waveforms from warp drive collapse

### v1

@@ 1 Introduction
Detections of gravitational waves by the LIGO Scientific-Virgo-Kagra Collaborations, e.g., [1, 2, 3] are paving the way to viewing extreme gravitational events across the entire visible universe. It is natural to ask what new signals originating from strongly distorted regions of spacetime could be seen in the future beyond the compact binary mergers already detected. To search for such signals and to correctly identify them in the measured data, it is important to understand their phenomenology and properties. Numerical relativity (NR) is one of the main tools for computing gravitational-wave signals for spacetimes in the highly dynamical strongly gravity regime in the absence of symmetries [4, 5, 6, 7]. NR simulations are time domain evolutions of a given initial state that allow the resulting gravitational-wave signals to be extracted in the asymptotic radiation zone, from which we can connect to our own observations from a given viewing angle and distance.
@@
Unlike black holes, which have a natural formation channel [8, 9] and considerable observational support [10, 11, 12], more exotic spacetimes such as wormholes or warp drives [13, 14] are considered science fiction because their formation and existence generate many potential paradoxes, and require matter that violates certain energy conditions. Nevertheless, warp drives can be described classically within the theory of general relativity by a metric first proposed by Miguel Alcubierre [15] and further developed in subsequent works, e.g., [16, 17, 18, 19].
@@
The principle idea behind a warp drive is that instead of exceeding the speed of light directly in a local reference frame, which would violate Lorentz invariance, a “warp bubble” could traverse distances faster than the speed of light (as measured by some distant observer) by contracting spacetime in front of it and expanding spacetime behind it [20]. The problems associated with such spacetimes are well documented  [21, 22, 23, 24, 13, 25]. One particular issue is the requirement for matter that violates the Null Energy Condition (NEC) [26, 27]  [^1] Recent works that have suggested positive energy warp drive solutions [28, 29, 30] have problems that are discussed in [31].. The requirement that warp drives violate the NEC may be considered a practical rather than fundamental barrier to their construction since NEC violation can be achieved by quantum effects and effective descriptions of modifications to gravity [32, 33], albeit subject to quantum inequality bounds [34, 35, 36] and other semiclassical considerations [37] that seem likely to prove problematic. Other issues with the warp drive metric include the potential for closed time-like curves [21, 23] and, from a more practical perspective, the difficulties for those in the ship in controlling and deactivating the bubble [38].
@@
This instability, whilst undesirable for the warp ship’s occupants, gives rise to the possibility of generating gravitational waves. The Alcubierre metric for a constant velocity warp ship would not generate gravitational waves. In addition, far away from the warp bubble, the spacetime is exactly Minkowskian – that is, it is actually flat, not just asymptotically flat. The warp bubble spacetime is therefore undetectable by its gravitational effects as it has no gravitational wave content and no measurable ADM mass [39]  [^2] It circumvents theorems about all asymptotically Minkowskian spacetimes being Minkowski [40, 41] by the presence of the negative energy contribution.. However, the collapse of the warp drive, or its acceleration or deceleration, should generate gravitational waves, and in this work, we study the first case by allowing the warp bubble to collapse. Physically, this could be related to a breakdown in the containment field that the post-warp civilisation (presumably) uses to support the warp bubble against collapse.
The prospect of a zero ADM mass spacetime radiating gravitational-wave energy immediately leads to an interesting physical question: Does the spacetime actually radiate energy, and if so, does the final spacetime have a negative ADM mass? In this work we consider the fluxes of energy both from the matter and the gravitational waves out of a spatial volume after the warp drive collapse. This study has connections to early work by Bondi [42], in which positive and negative matter restricted to linear motion was shown to attract and repel in a way that would accelerate them together.
For our study, we restrict ourselves to a limited and simplified scenario and sub-light speeds as a proof of principle [^3] Our set-up appears to work for speeds of order $v\lesssim 1$ but requires much higher resolutions for larger speeds. It would be better to use a code adapted to the cylindrical symmetry of the problem to study such cases.. The initial bubble is described by the original Alcubierre metric with a fixed wall thickness. We develop a formalism to describe the warp fluid and its evolution and vary its initial velocity at the point of collapse (which is related to the amplitude of the warp bubble). We neglect the existence of the warp ship itself [27], which is assumed to be negligible compared to the bubble.
@@
Figure 1: The evolution in time of the trace of $K_{ij}$, $K=\gamma^{ij}K_{ij}$. This quantity initially provides a schematic representation of the warp drive metric, with positive values for the contracting part of the space ahead of the ship and negative values for the expanding part behind. We observe the destabilisation of these regions, which initially collapse inwards and then dissipate to infinity. Movie at https://youtu.be/6aqkxooXAc4?si=NhremdNv5_F3sTKL.
@@ 2 Formalism for the warp drive simulation
as an initial value problem, using the $3+1$-decomposed form and setting $G=c=1$. This means that we define spacelike hypersurfaces that foliate the four-dimensional manifold, specify initial data on a particular hypersurface, and evolve it forward in our chosen time coordinate using the method of lines. (See e.g. [43, 44, 5] for detailed discussions.) The $3+1$-decomposed metric in coordinates adapted to the slicing takes the form
@@ 2.1 Initial data for the metric
Although, as discussed above, many alternatives with improved physical properties have been discussed in the literature, we will consider in this work the original Alcubierre metric [15] for which
@@ 2.3 Metric evolution
Figure 2: The evolution in time of the real part of the Weyl scalar $\Psi_{4}$ for the case of $v=0.1$. This quantity provides a measure of the spacetime curvature and, in the far region, can be identified with the gravitational wave content of the spacetime. In the last two panels, we see a burst of gravitational-wave radiation leaving the collapsed remnant of the warp bubble. Movie at https://youtu.be/h-BieIMO9Xw?si=vVFQf0k9nZiQGQkz.
For the metric evolution, we use the standard CCZ4 formulation of [46] and the moving puncture gauge [47, 48, 49, 50] within the numerical-relativity code grchombo [51, 52, 53], which uses the method of lines, with an RK4 time integration and 4th order finite difference stencils for calculating spatial gradients.
We note that for the gauge evolution we use the shock-avoiding Bona-Masso type gauge slicing originally proposed by Alcubierre [54, 55], and recently advocated for by Baumgarte and Hilditch [56] in the context of critical collapse simulations. We find that the shock-avoiding slicing gives much better stability and, more importantly, allows us to extract coherent gravitational waveforms in the wave zone, whereas in the standard gauge, we obtained large unphysical gauge effects that hindered us in extracting the physical gravitational-wave signal. Because this gauge has a tendency to push the extrinsic curvature to zero, one should not over-interpret the evolution shown in Fig. 1. It quickly decays to zero in the simulation before the matter and gravitational-wave fluxes have begun to leave the volume.
Better quantities to track the evolution of the metric curvature are the Weyl scalars, in particular $\Psi_{4}$. We study the tensorial gravitational-wave modes emitted by the warp drive by extracting the Newman-Penrose scalar $\Psi_{4}$ with the tetrads proposed by [57], projected into spin-weight 2-2 spherical harmonics, $\psi_{lm}=\oint_{S^{2}}\Psi_{4}|_{r=r_{\mathrm{ex}}}\left[{}_{-2}\bar{Y}^{lm}\right]\,\rmd\Omega$, where $\rmd\Omega=\sin\theta\,\rmd\theta\,\rmd\varphi$ is the area element on the $S^{2}$ unit sphere. The evolution of the real part of $\Psi_{4}$ is illustrated in Fig. 2.
@@ 2.4 Matter evolution
Figure 3: The evolution in time of the matter energy density $\rho$ for the case of $v=0.1$. We see that a ring of positive energy forms within the initially negative energy density bubble and propels it outwards, which is reminiscent of [42]. There are several waves of alternating positive and negative energies. In the final panel, we see that the remnant appears stable, and compared to Fig. 2, we see that the matter waves propagate at roughly the same speed as the gravitational waves. Movie at https://youtu.be/RGRB5Y0cl7w?si=e6zln3_ijxRDNd44.
@@
The evolution equation for $S_{ij}$ is, in principle, free - once we specify the equation of state of the matter, it would be fully determined. We base our choice on the observation that the initial warp drive spacetime has an average pressure of [31]
@@
This construction assumes that there is some transition after the warp field breakdown that reverts the warp fluid to a (phantom) perfect fluid with a stiff equation of state and a velocity that matches the normal observers, such that the divergence term in Eq. (19) and the anisotropic stresses decay away. Further investigations are needed to understand the influence of the equation of state of the warp field on the gravitational-wave signal, but it appears that the metric behaviour mostly dominates the signal. However, the chosen equation of state will strongly affect the matter fluxes, which may change the conclusions in Sec. 3.
@@ 2.5 Numerical setup
The size of the domain is L=168[R]L=168~[R] with $N=224$ cells on the coarsest grid in the direction of travel of the warp ship (the $z$ axis). In the $x$ and $y$ directions, we use the symmetry of the problem to halve the simulation domain so that we only evolve a quarter of the full domain, with appropriate parities at the reflection boundaries. In the nonsymmetric directions, we use Sommerfeld boundary conditions; cf. [5]. We employ 6 levels of 2:1 refinement, such that the finest grid has a width of $2.625[R]$ and the finest grid spacing is $dx\sim 0.0117[R]$.
For all runs, the size of the warp ship in code units is set to be $R=1$, and the value that determines the wall thickness $\sigma$ is set to be $8$. The velocity of the warp ship is varied between 0.1 and 0.2. We have found that we can stably evolve higher velocities up to 0.5, but good quality results require very high resolutions, which would be better achieved with a code adapted to the cylindrical symmetry. Above $v=1$, we expect to encounter pathologies related to the ship’s faster-than-light speed, but again, we would require higher resolutions and more computational resources to confirm this.
Details of the checks we have done on convergence and verification of the matter evolution are given in A.
@@ 3.1 Gravitational waves
A visualisation of the evolution of the gravitational-wave content is shown in Fig. 2. We plot the extracted Weyl scalar $\Psi_{4}$ for the main $m=0$ modes in Fig. 4. The signal comes as a burst, initially having no gravitational wave content, followed by an oscillatory period with a characteristic frequency of order $1/[R]$. Overall, the signal is very distinct from the typical compact binary coalescences observed by gravitational-wave detectors and more similar to events like the collapse of an unstable neutron star or the head-on collision of two black holes [58, 59]. However, since the remnant is not a black hole, there is no characteristic black hole ringdown in the signal, and there is evidence of a longer lower frequency tail, particularly for higher velocities.
Figure 4: Modes of the Newman-Penrose scalar $\Psi_{4}$ multiplied by the coordinate radius for different values of the warp speed $v$. Given the symmetry of the problem, the $\neq 0 m$odes are zero, so we only show the modes with $l=2,3,4$.
Integrating $\Psi_{4}$, we find that for a warp speed of $v=0.1$, the magnitude of the strain times the radius of extraction is of order $10^{-2}[R]$. Thus, for a 1 km-sized warp ship, the strain of the signal at a distance of 1 Mpc would be $10^{-21}$. However, its frequency ($f\sim 1/[R]$) equates to $f\sim$ 300 kHz, which is outside the range of current ground based detectors. Proposals for higher frequency detectors have been made [60], so in future one may be able to put bounds on the existence of such signals. Depending on the size of the warp bubble, which determines the frequency of the signal, as well as the distance to the source, these signatures may be detected through a coherent search or burst search combining information from multiple gravitational-wave detectors, e.g., [61, 62, 63]. A more in depth analysis, like the one of [64], would be needed to give accurate bounds.
Figure 5: A comparison of the matter and gravitational-wave energy fluxes out of sphere at a coordinate radius of $r=14[R]$ around the warp bubble versus time for different $v$. The left panel shows the integrated GW fluxes in the 20 mode, whereas the right shows the integrated matter fluxes. The net gravitational-wave flux is always positive, whereas the net matter flux oscillates as waves of positive and negative energy leave the volume.
Figure 6: Zoom in on the start and end periods of the integrated energy fluxes in the matter and gravitational waves as in Fig. 5 for the case $v=0.1$. We see that the flux begins in the matter sector, with alternating waves of positive and negative energy. The gravitational-wave flux starts a little later but continues throughout the periods in which the net matter flux is both positive and negative. The second panel shows that the overall matter energy loss is negative (i.e. the energy of matter in the volume increases), whereas the gravitational-wave flux is overall positive as expected but much smaller. The final ADM mass of the volume is therefore larger than its initial value (which was zero).

### v2

@@ 1 Introduction
Detections of gravitational waves by the LIGO Scientific-Virgo-Kagra Collaborations (e.g., Abbott et al., 2018, 2021, 2023) are paving the way to viewing extreme gravitational events across the entire visible universe. It is natural to ask what new signals originating from strongly distorted regions of spacetime could be seen in the future beyond the compact binary mergers already detected. To search for such signals and to correctly identify them in the measured data, it is important to understand their phenomenology and properties. Numerical relativity (NR) is one of the main tools for computing gravitational-wave signals for spacetimes in the highly dynamical strongly gravity regime in the absence of symmetries (Gourgoulhon, 2007; Alcubierre, 2008; Baumgarte and Shapiro, 2010, 2021). NR simulations are time domain evolutions of a given initial state that allow the resulting gravitational-wave signals to be extracted in the asymptotic radiation zone, from which we can connect to our own observations from a given viewing angle and distance.
@@
Unlike black holes, which have a natural formation channel (Chandrasekhar, 1931; Penrose, 1965) and considerable observational support (Akiyama et al., 2019; Abbott et al., 2016; Ghez et al., 2008), more exotic spacetimes such as wormholes or warp drives (Lobo, 2007; Santiago et al., 2021) are considered science fiction because their formation and existence generate many potential paradoxes, and require matter that violates certain energy conditions. Nevertheless, warp drives can be described classically within the theory of general relativity by a metric first proposed by Miguel Alcubierre (Alcubierre, 1994) and further developed in subsequent works (e.g., Natário, 2002; Van Den Broeck, 1999; Krasnikov, 1998; Everett and Roman, 1997).
@@
The principle idea behind a warp drive is that instead of exceeding the speed of light directly in a local reference frame, which would violate Lorentz invariance, a “warp bubble” could traverse distances faster than the speed of light (as measured by some distant observer) by contracting spacetime in front of it and expanding spacetime behind it (Alcubierre and Lobo, 2017). The problems associated with such spacetimes are well documented  (Everett, 1996; Hiscock, 1997; Visser et al., 2000; Loup et al., 2001; Lobo, 2007; Obousy and Cleaver, 2008). One particular issue is the requirement for matter that violates the Null Energy Condition (NEC) (Olum, 1998; Lobo and Visser, 2004)  [^1] Recent works that have suggested positive energy warp drive solutions (Lentz, 2021; Bobrick and Martire, 2021; Fell and Heisenberg, 2021) have problems that are discussed in (Santiago et al., 2022).. The requirement that warp drives violate the NEC may be considered a practical rather than fundamental barrier to their construction since NEC violation can be achieved by quantum effects and effective descriptions of modifications to gravity (Barcelo and Visser, 2002; Kontou and Sanders, 2020), albeit subject to quantum inequality bounds (Pfenning and Ford, 1997; Fewster and Teo, 1999; Krasnikov, 2003) and other semiclassical considerations (Finazzi et al., 2009) that seem likely to prove problematic. Other issues with the warp drive metric include the potential for closed time-like curves (Everett, 1996; Visser et al., 2000) and, from a more practical perspective, the difficulties for those in the ship in controlling and deactivating the bubble (Alcubierre, 2017).
@@
This instability, whilst undesirable for the warp ship’s occupants, gives rise to the possibility of generating gravitational waves. The Alcubierre metric for a constant velocity warp ship would not generate gravitational waves. In addition, far away from the warp bubble, the spacetime is exactly Minkowskian – that is, it is actually flat, not just asymptotically flat. The warp bubble spacetime is therefore undetectable by its gravitational effects as it has no gravitational wave content and no measurable ADM mass (Schuster et al., 2023)  [^2] It circumvents theorems about all asymptotically Minkowskian spacetimes being Minkowski (Witten, 1981; Schon and Yau, 1979) by the presence of the negative energy contribution. Note that the spacetime at spatial infinity, where the ADM mass is defined, is causally disconnected from the spacetime of the warp bubble, therefore the true ADM mass should be conserved and remain at zero. When we refer to the mass of the spacetime volume here, after a period of dynamical evolution, we mean a quasi-local analogy obtained by integrating over the surface of a finite volume some distance away from the collapsing bubble, after that volume has reached a stationary state with no net fluxes in or out. (although other signatures are possible, see e.g. (Marzlin and Kinach, 2022)). However, the collapse of the warp drive, or its acceleration or deceleration, should generate gravitational waves, and in this work, we study the first case by allowing the warp bubble to collapse. Physically, this could be related to a breakdown in the containment field that the post-warp civilisation (presumably) uses to support the warp bubble against collapse.
The prospect of a zero ADM mass spacetime radiating gravitational-wave energy immediately leads to an interesting physical question: Does the spacetime actually radiate energy, and if so, once it settles down to a steady state, does the final spacetime have a negative mass? In this work we consider the fluxes of energy both from the matter and the gravitational waves out of a spatial volume after the warp drive collapse. This study has connections to early work by Bondi (Bondi, 1957), in which positive and negative matter restricted to linear motion was shown to attract and repel in a way that would accelerate them together.
For our study, we restrict ourselves to a limited and simplified scenario and sub-light speeds as a proof of principle. This is also a first step towards studying what, if any, pathologies develop as we reach $v=c$. We would also emphasise that the warp drive is interesting since it uses gravitational distortion of spacetime to propel the ship forward, rather than a usual kind of fuel/reaction system. This means that its detection would signal the use of NEC violating drives, even in sub-light-speed cases.
The initial bubble is described by the original Alcubierre metric with a fixed wall thickness. We develop a formalism to describe the warp fluid and its evolution and vary its initial velocity at the point of collapse (which is related to the amplitude of the warp bubble). We neglect the existence of the warp ship itself (Lobo and Visser, 2004), which is assumed to be negligible compared to the bubble.
@@
Figure 1: The evolution in time of the trace of $K_{ij}$, $K=\gamma^{ij}K_{ij}$. This quantity initially provides a schematic representation of the warp drive metric, with positive values for the contracting part of the space ahead of the ship and negative values for the expanding part behind. We observe the destabilisation of these regions, which initially collapse inwards and then dissipate to infinity. Movie at https://youtu.be/rZsre-y4hlM?feature=shared.
@@ 2 Formalism for the warp drive simulation
as an initial value problem, using the $3+1$-decomposed form and setting $G=c=1$. This means that we define spacelike hypersurfaces that foliate the four-dimensional manifold, specify initial data on a particular hypersurface, and evolve it forward in our chosen time coordinate using the method of lines. (See the standard numerical relativity texts  (Arnotwitt et al., 1960; York, 1979; Alcubierre, 2008) for detailed discussions.) The $3+1$-decomposed metric in coordinates adapted to the slicing takes the form
@@ 2.1 Initial data for the metric
Although, as discussed above, many alternatives with improved physical properties have been discussed in the literature, we will consider in this work the original Alcubierre metric (Alcubierre, 1994) for which
@@ 2.3 Metric evolution
Figure 2: The evolution in time of the real part of the Weyl scalar $\Psi_{4}$ for the case of $v=0.1$. This quantity provides a measure of the spacetime curvature and, in the far region, can be identified with the gravitational wave content of the spacetime. In the last two panels, we see a burst of gravitational-wave radiation leaving the collapsed remnant of the warp bubble. Movie at https://youtu.be/rZsre-y4hlM?feature=shared.
For the metric evolution, we use the standard CCZ4 formulation of (Alic et al., 2012) and the moving puncture gauge (Bona et al., 1995; Baker et al., 2006; Campanelli et al., 2006; van Meter et al., 2006) within the numerical-relativity code grchombo (Andrade et al., 2021; Radia et al., 2022; Clough et al., 2015), which uses the method of lines, with an RK4 time integration and 4th order finite difference stencils for calculating spatial gradients.
We note that for the gauge evolution we use the shock-avoiding Bona-Masso type gauge slicing originally proposed by Alcubierre (Alcubierre, 1997, 2003), and recently advocated for by Baumgarte and Hilditch (Baumgarte and Hilditch, 2022) in the context of critical collapse simulations. We find that the shock-avoiding slicing gives much better stability and, more importantly, allows us to extract coherent gravitational waveforms in the wave zone, whereas in the standard gauge, we obtained large unphysical gauge effects that hindered us in extracting the physical gravitational-wave signal. Because this gauge has a tendency to push the extrinsic curvature to zero, one should not over-interpret the evolution shown in Fig. 1. It quickly decays to zero in the simulation before the matter and gravitational-wave fluxes have begun to leave the volume.
[…]
@@ 2.4 Matter evolution
Figure 3: The evolution in time of the matter energy density $\rho$ for the case of $v=0.1$. We see that a ring of positive energy forms within the initially negative energy density bubble and propels it outwards, which is reminiscent of (Bondi, 1957). There are several waves of alternating positive and negative energies. In the final panel, we see that the remnant appears stable, and compared to Fig. 2, we see that the matter waves propagate at roughly the same speed as the gravitational waves. Movie at https://youtu.be/rZsre-y4hlM?feature=shared.
@@
The evolution equation for $S_{ij}$ is, in principle, free - once we specify the equation of state of the matter, it would be fully determined. We base our choice on the observation that the initial warp drive spacetime has an average pressure of (Santiago et al., 2022)
@@
This construction assumes that there is some transition after the warp field breakdown that reverts the warp fluid to a perfect fluid with a stiff equation of state and a velocity that matches the normal observers, such that the divergence term in Eq. (19) and the anisotropic stresses decay away. Further investigations are needed to understand the influence of the equation of state of the warp field on the gravitational-wave signal, but it appears that the metric behaviour mostly dominates the signal. However, the chosen equation of state will strongly affect the matter fluxes, which may change the conclusions in Sec. 3.
@@ 2.5 Numerical setup
The size of the domain is L=168[R]L=168~[R] with $N=224$ cells on the coarsest grid in the direction of travel of the warp ship (the $z$ axis). In the $x$ and $y$ directions, we use the symmetry of the problem to halve the simulation domain so that we only evolve a quarter of the full domain, with appropriate parities at the reflection boundaries. In the nonsymmetric directions, we use Sommerfeld boundary conditions  (Radia et al., 2022). We employ 6 levels of 2:1 refinement, such that the finest grid has a width of $2.625[R]$ and the finest grid spacing is $dx\sim 0.0117[R]$.
For all runs, the size of the warp ship in code units is set to be $R=1$, and the value that determines the wall thickness $\sigma$ is set to be $8$. The velocity of the warp ship is varied between 0.1 and 0.2. We have found that we can stably evolve higher velocities up to 0.5, but good quality results require very high resolutions, which would be better achieved with a code adapted to the axisymmetry. This work can be viewed as a step towards the study of the physicality of the faster-than-light case, where there are good reasons to suppose that the spacetime should be pathological in some way. By showing that numerical evolutions can be made to work at sub-light speeds, this gives us hope of being able to trace a point of breakdown near light speed.
The waves are extracted at a radius of $14[R]$. This is much smaller than values used in binary extraction, but the system here is very different and so the comparison is not appropriate. In particular, the spacetime is already flat outside the original bubble wall at $R$. The total integrated mass of the matter is also initially of order (minus) $M=\int\rho dV\sim 0.01R$, and not $M\sim R$ as in, say, a neutron star. Since the frequency of the waves is high relative to the grid spacing, we need to extract further in to obtain a good resolution of the frequencies using the adaptive mesh refinement (AMR) hierarchy. We have checked the consistency of the waves at different resolutions, and details of the checks we have done on convergence of the constraints and verification of the matter evolution are given in Appendix A.
@@ 3.1 Gravitational waves
A visualisation of the evolution of the gravitational-wave content is shown in Fig. 2. We plot the extracted Weyl scalar $\Psi_{4}$ for the main $m=0$ modes in Fig. 4. The signal comes as a burst, initially having no gravitational wave content, followed by an oscillatory period with a characteristic frequency of order $1/[R]$. Overall, the signal is very distinct from the typical compact binary coalescences observed by gravitational-wave detectors and more similar to events like the collapse of an unstable neutron star or the head-on collision of two black holes (Giacomazzo et al., 2011; Sperhake et al., 2005). However, since the remnant is not a black hole, there is no characteristic black hole ringdown in the signal, and there is evidence of a longer lower frequency tail, particularly for higher velocities.
Integrating $\Psi_{4}$, we find that for a warp speed of $v=0.1$, the magnitude of the strain times the radius of extraction is of order $10^{-2}[R]$. Thus, for a 1 km-sized warp ship, the strain of the signal at a distance of 1 Mpc would be $10^{-21}$. However, its frequency ($f\sim 1/[R]$) equates to $f\sim$ 300 kHz, which is outside the range of current ground based detectors. Proposals for higher frequency detectors have been made (Aggarwal et al., 2021), so in future one may be able to put bounds on the existence of such signals. Depending on the size of the warp bubble, which determines the frequency of the signal, as well as the distance to the source, these signatures may be detected through a coherent search or burst search combining information from multiple gravitational-wave detectors (e.g., Sutton et al., 2010; Cornish and Littenberg, 2015; Klimenko et al., 2016). A more in depth analysis, like the one of Kuwahara and Cannon (2023), would be needed to give accurate bounds.
We note that since the matter and gravitational waves propagate at the same speed, the gravitational waves are never in an exact vacuum. We use the appropriate calculation of the Weyl scalars to account for this, but it means that the two components may continue to interact and that the gravitational waves will not fall off exactly as $1/r$ (although the difference is small over the domain we study). This is specific to the case where the matter propagates at the speed of light, whereas in a more realistic matter model we would expect the matter to be moving slower and the gravitational waves to propagate ahead. This is another point that could be investigated further in future work.
Figure 4: Modes of the Newman-Penrose scalar $\Psi_{4}$ multiplied by the coordinate radius for different values of the warp speed $v$. Given the symmetry of the problem, the $\neq 0 m$odes are zero, so we only show the modes with $l=2,3,4$. We observe that the signal has a frequency set by the size of the warp bubble $f\sim 1/[R]$, which for a 1 km warp bubble would be of the order $f\sim 300$ kHz.